# Breaking the Dimensional Barrier: Dynamic Portfolio Choice with Parameter Uncertainty via Pontryagin Projection

### v2

@@ E.5 Engineering notes and stabilizers
    with $c\in(0,1)$ and $\varepsilon>0$.
  • Initialization and scale control in high dimensions. To avoid early-time numerical blow-ups (often through quadratic variation terms of the form $\pi^{\top}\Sigma\pi$), initialize the policy output near zero and/or scale the output by $d^{-1/2}$. As a last-resort safety net, a mild log-wealth clamp can prevent overflow, but it should be used conservatively and monitored, since frequent clamping may distort higher-order sensitivities.
Appendix F Coupling stage 1 and stage 2: residual projection and interactive distillation
We keep the ex–ante objective (48) and the $\theta$-blind deployability constraint throughout. Stage 2 is not a separate optimization problem: it is a (warm-started) post-processing map that reuses the current stage 1 policy as a warm-up control, estimates the projection ingredients under this warm-up policy, and then applies the $q$-aggregated Pontryagin projection as a deterministic transformation of the policy. Concretely, stage 2 takes $\pi^{\mathrm{warm}}$ as input, runs BPTT/Monte Carlo under frozen $\theta\sim q$ to estimate the adjoint blocks required for the Hamiltonian gradient, aggregates them across $\theta$, and outputs a single deployable $\theta$-blind rule $\widehat{\pi}^{\mathrm{agg,mix}}$ (Section 3.2).
This appendix records two couplings between the two stages, each with a distinct role:
  • Residual (control-variate) projection. We compute the same projected rule in a residual form around the warm-up policy. This is algebraically equivalent to the direct projection, but typically reduces Monte Carlo variance and improves numerical stability in high dimensions.
  • Interactive distillation (amortized projection). We use the projected rule as a teacher signal at intermittent refresh times. Beyond accelerating/stabilizing training, distillation amortizes the cost of projection: stage 2 can be accurate but Monte-Carlo intensive, whereas a distilled student policy approximates the projected rule with a single forward pass at stage 1 inference cost.
F.1 Control-variate (residual) form of the projected rule
@@
which is an algebraic rewriting of (59) (hence it does not change the target).
Practical value (variance reduction and stability). When the warm-up policy is already close to a projected fixed point on the working domain, the residual $\widehat{r}_{\mathrm{FOC}}$ tends to be small. Moreover, the ingredients entering $\widehat{A}_{t}\pi_{\varphi^{\mathrm{warm}}}$ and $\widehat{G}_{t}^{\mathrm{mix}}$ are computed from the same Monte Carlo pool and often partially cancel, improving concentration of $\widehat{r}_{\mathrm{FOC}}$. In implementation, we combine (104) with the inversion stabilizers of Appendix E.5 (e.g. ridge regularization and diagnostic-based fallback to $\pi_{\varphi^{\mathrm{warm}}}$ on unreliable states).
F.2 Interactive distillation: projection-guided training and amortized deployment
Let $\pi_{\varphi}$ be the trainable stage 1 policy network. At intermittent refresh times, we freeze a lagged copy $\pi_{\varphi^{-}}$ and run stage 2 under $\pi_{\varphi^{-}}$ to construct a $q$-aggregated projected teacher. This coupling serves two purposes:
  • Training aid. Projection-guided targets can stabilize and accelerate stage 1 optimization, especially once the warm-up policy enters a locally stable regime.
  • Amortization. Stage 2 projection can be accurate but Monte-Carlo intensive; distillation compresses it into a single forward pass for deployment.
Teacher definition (computed by stage 2, treated as fixed).
In residual form (104), the teacher is the $\theta$-blind map
@@ F.2 Interactive distillation: projection-guided training and amortized deployment
where $\widehat{r}_{\mathrm{FOC}}^{\varphi^{-}}:=\widehat{A}_{t}^{\varphi^{-}}\pi_{\varphi^{-}}+\widehat{G}_{t}^{\mathrm{mix},\varphi^{-}}$ is computed using the mixed-moment $q$-aggregation under the lagged policy. In practice, the inversion in (105) is implemented with the same stabilizers as in Section 3.2 (ridge/diagnostics/residual form); for notational simplicity we keep the formal inverse.
Student objective (hybrid PG–DPO + distillation loss).
We train $\pi_{\varphi}$ by combining the original ex–ante objective with a proximity term to the teacher on the working domain:
$\max_{\varphi}\;J(\varphi)\;-\;\lambda\,\mathbb{E}_{(t,x,y)\sim\mu}\Big[\big\|\pi_{\varphi}(t,x,y)-\mathrm{stopgrad}\big(\widehat{\pi}^{\mathrm{agg,mix}}_{\varphi^{-}}(t,x,y)\big)\big\|^{2}\Big],$ (106)
where $\mu$ is the working-domain sampling measure and $\lambda\geq 0$ controls the strength of projection guidance. The operator $\mathrm{stopgrad}(\cdot)$ indicates that gradients are not propagated through stage 2: once computed from $\pi_{\varphi^{-}}$, the teacher is treated as fixed.
Refresh schedule, annealing, and diagnostic gating (implementation).
In practice, $\varphi^{-}$ and $\widehat{\pi}^{\mathrm{agg,mix}}_{\varphi^{-}}$ are refreshed on a slower timescale than the stage 1 gradient steps. A typical and robust configuration is:
  • Two-time-scale refresh. Hold $\varphi^{-}$ fixed for $K_{\mathrm{refresh}}$ stage 1 updates of $\varphi$ under (106), then set $\varphi^{-}\leftarrow\varphi$ and recompute the teacher via stage 2.
  • Anneal $\lambda$. Start with $\lambda=0$ (pure PG–DPO) and increase $\lambda$ only after basic projection checks on the working domain indicate that the stage 2 map is reliable. Concretely, we anneal $\lambda$ toward a target value $\lambda_{\max}$ once the warm-up residual $\|\widehat{r}^{\mathrm{warm}}_{\mathrm{FOC,mix}}\|_{L^{2}(\mu)}$ and curvature diagnostics remain stable over several refresh cycles.
  • Adaptive teacher selection (gating). To avoid injecting noisy teacher targets early in training or on pathological regions of the domain, we optionally apply the teacher loss only on states where projection diagnostics certify reliability (e.g. bounded condition number of $\widehat{A}_{t}$, eigenvalue/denominator lower bounds, and moderate residual magnitude), falling back to pure PG–DPO elsewhere.
Implementation details, including the precise diagnostics and stabilizers, are deferred to Appendix E.
Deployment. After training, the deployed controller is simply the student network $\pi_{\varphi}$, which approximates the stage 2 projected rule with a single forward pass, without repeatedly running BPTT/Monte Carlo projection at inference time.
Appendix G Stage 2 projection diagnostics
We report Stage 2 diagnostic statistics as a visual supplement to Section 4.2. Each figure summarizes the same tail-median protocol and layout; see captions for definitions and interpretation.
@@ Appendix G Stage 2 projection diagnostics
Appendix H Supplementary decomposition diagnostics for Section 5
Tables 23 report Stage 2 decomposition diagnostics at $t=0$.
@@ Appendix H Supplementary decomposition diagnostics for Section 5
$10^{-3}$ Stage 1+Stage 2 (Basic) $7.17\text{\times}{10}^{-06}$ $1.51\text{\times}{10}^{-05}$ $1.60\text{\times}{10}^{-05}$ $1.23\text{\times}{10}^{-05}$
Stage 1+Stage 2 (Distill.) $5.20\text{\times}{10}^{-06}$ $9.81\text{\times}{10}^{-06}$ $1.64\text{\times}{10}^{-05}$ $1.62\text{\times}{10}^{-05}$
$10^{-2}$ Stage 1+Stage 2 (Basic) $6.21\text{\times}{10}^{-06}$ $1.38\text{\times}{10}^{-05}$ $1.63\text{\times}{10}^{-05}$ $1.42\text{\times}{10}^{-05}$
Stage 1+Stage 2 (Distill.) $7.13\text{\times}{10}^{-06}$ $7.13\text{\times}{10}^{-06}$ $1.62\text{\times}{10}^{-05}$ $1.76\text{\times}{10}^{-05}$
$10^{-1}$ Stage 1+Stage 2 (Basic) $8.18\text{\times}{10}^{-06}$ $1.41\text{\times}{10}^{-05}$ $3.82\text{\times}{10}^{-05}$ $2.42\text{\times}{10}^{-05}$
Stage 1+Stage 2 (Distill.) $6.72\text{\times}{10}^{-06}$ $9.21\text{\times}{10}^{-06}$ $3.67\text{\times}{10}^{-05}$ $3.16\text{\times}{10}^{-05}$
@@
$10^{-3}$ Stage 1+Stage 2 (Basic) $1.10\text{\times}{10}^{-05}$ $1.41\text{\times}{10}^{-05}$ $70\text{\times}{10}^{-05} 1.1$$18\text{\times}{10}^{-05}$
Stage 1+Stage 2 (Distill.) $7.71\text{\times}{10}^{-06}$ $5.94\text{\times}{10}^{-06}$ $1.84\text{\times}{10}^{-05}$ $1.62\text{\times}{10}^{-05}$
$10^{-2}$ Stage 1+Stage 2 (Basic) $7.90\text{\times}{10}^{-06}$ $2.02\text{\times}{10}^{-05}$ $1.46\text{\times}{10}^{-05}$ $1.20\text{\times}{10}^{-05}$
Stage 1+Stage 2 (Distill.) $6.00\text{\times}{10}^{-06}$ $1.71\text{\times}{10}^{-05}$ $2.21\text{\times}{10}^{-05}$ $1.43\text{\times}{10}^{-05}$
$10^{-1}$ Stage 1+Stage 2 (Basic) $1.24\text{\times}{10}^{-05}$ $1.93\text{\times}{10}^{-04}$ $5.77\text{\times}{10}^{-05}$ $3.14\text{\times}{10}^{-05}$
Stage 1+Stage 2 (Distill.) $1.20\text{\times}{10}^{-05}$ $1.90\text{\times}{10}^{-04}$ $7.40\text{\times}{10}^{-05}$ $2.47\text{\times}{10}^{-05}$
Table 2: Myopic-component RMSE at $t=0$ (tail medians).
@@
$10^{-3}$ Stage 1+Stage 2 (Basic) $0.994$ $0.988$ $0.991$ $0.990$
Stage 1+Stage 2 (Distill.) $0.995$ $0.986$ $0.990$ $0.987$
$10^{-2}$ Stage 1+Stage 2 (Basic) $0.993$ $0.989$ $0.992$ $0.988$
Stage 1+Stage 2 (Distill.) $0.992$ $0.994$ $0.990$ $0.987$
$10^{-1}$ Stage 1+Stage 2 (Basic) $0.988$ $0.990$ $0.936$ $0.932$
Stage 1+Stage 2 (Distill.) $0.996$ $0.990$ $0.949$ $0.922$
@@
$10^{-3}$ Stage 1+Stage 2 (Basic) $0.988$ $0.988$ $0.993$ $0.990$
Stage 1+Stage 2 (Distill.) $0.994$ $0.995$ $0.990$ $0.987$
$10^{-2}$ Stage 1+Stage 2 (Basic) $0.994$ $0.976$ $0.992$ $0.988$
Stage 1+Stage 2 (Distill.) $0.994$ $0.980$ $0.988$ $0.989$
$10^{-1}$ Stage 1+Stage 2 (Basic) $0.990$ $0.005$ $0.668$ $0.851$
Stage 1+Stage 2 (Distill.) $0.992$ 0.009-0.009 $0.642$ $0.871$
Table 3: Hedging-direction cosine similarity at $t=0$ (tail medians). Higher is better; negative indicates direction reversal.

### v3

@@ F.5 Engineering notes and stabilizers
    with $c\in(0,1)$, $\varepsilon>0$.
  • Initialization/scale control in high dimensions. Initialize policy outputs near zero and/or scale outputs by $d^{-1/2}$ to avoid early-time blow-ups (e.g. via $\pi^{\top}\Sigma\pi$).
Appendix G Coupling stage 1 and stage 2: residual projection and interactive distillation
We keep the ex–ante objective (42) and the $\theta$-blind deployability constraint throughout. Stage 2 is not a separate optimization problem: it is a (warm-started) post-processing map that reuses the current stage 1 policy as a warm-up control, estimates projection ingredients under this warm-up policy, and then applies a $q$-aggregated Pontryagin projection as a deterministic transformation.
Concretely, stage 2 takes $\pi^{\mathrm{warm}}$ as input, runs BPTT/Monte Carlo under frozen $\theta\sim q$ to estimate the adjoint blocks required for the Hamiltonian gradient, aggregates them across $\theta$, and outputs a single deployable $\theta$-blind rule $\widehat{\pi}^{\mathrm{agg,mix}}$ (Section 3.2).
This appendix records two couplings between the two stages:
  • Residual (control-variate) projection. The projected rule is evaluated in a residual form around the warm-up policy. This is algebraically equivalent to the direct solve, but can reduce Monte Carlo variance and improve numerical stability.
  • Interactive distillation (amortized projection). The projected rule is used as a teacher signal at intermittent refresh times. Distillation amortizes the cost of projection by compressing projected controls into the policy network.
Stage 1: DPO (Section 3.1)
Goal. Maximize $J(\varphi)=\mathbb{E}[U(X_{T}^{\pi_{\varphi},\theta})]$ with a deployable $\theta$-blind policy.
Loop. Sample $(t_{0},x_{0},y_{0})\sim\nu$ and latent $\theta\sim q$ inside the simulator; simulate Euler rollouts; compute $U(X_{T})$; BPTT for $\nabla_{\varphi}U(X_{T})$; update $\varphi$ (e.g. Adam). Output. Warm-up checkpoint $\varphi^{\mathrm{warm}}$.
Stage 2: $q$-aggregated Pontryagin projection under latent $\theta$ (Section 3.2)
Input. $\pi_{\varphi^{\mathrm{warm}}}$, working-domain sampler $\mu$ on $\mathcal{D}$, budgets $(M_{\mathrm{MC}},M_{\theta})$.
Adjoint blocks. Simulate under $\pi_{\varphi^{\mathrm{warm}}}$ and compute $\widehat{p}_{t}^{\theta}(z),\widehat{p}_{x,t}^{\theta}(z),\widehat{p}_{y,t}^{\theta}(z)$ by autodiff/BPTT. Projection. Aggregate $\widehat{A}_{t}$ and $\widehat{g}_{t}^{\mathrm{mix}}$ via (51)–(52); output $\widehat{\pi}^{\mathrm{agg,mix}}(z)=-(\widehat{A}_{t}(z))^{-1}\widehat{g}_{t}^{\mathrm{mix}}(z)$ (53). Stabilizers. Use ridge/diagnostics/residual form as needed (Appendix F.5). Output. Accurate but Monte-Carlo intensive projected rule.
Coupling (Appendix G)
(G.1) Residual projection.
Compute $\widehat{r}_{\mathrm{FOC}}(z)=\widehat{A}_{t}(z)\pi_{\varphi^{\mathrm{warm}}}(z)+\widehat{g}_{t}^{\mathrm{mix}}(z)$ and evaluate $\widehat{\pi}^{\mathrm{agg,mix}}(z)=\pi_{\varphi^{\mathrm{warm}}}(z)-\widehat{A}_{t}(z)^{-1}\widehat{r}_{\mathrm{FOC}}(z)$ (110).
(G.2) Interactive distillation (amortized projection).
Freeze a lagged copy $\varphi^{-}$, compute a teacher via stage 2, and train the student with the mixed objective (112).
Refresh teacher on a slower timescale and optionally gate teacher loss by projection diagnostics.
[…]
Figure 4: Pipeline of Section 3: stage 1 learning, stage 2 $q$-aggregated projection, and the coupling mechanisms in Appendix G. Distillation stabilizes training and amortizes projection by compressing projected controls into the policy network.
G.1 Control-variate (residual) form of the projected rule
@@
which is algebraically equivalent to (53) (hence it does not change the target).
Practical value (variance reduction and stability). When the warm-up policy is already close to a projected fixed point on the working domain, the residual $\widehat{r}_{\mathrm{FOC}}$ tends to be small. Moreover, the ingredients entering $\widehat{A}_{t}\pi_{\varphi^{\mathrm{warm}}}$ and $\widehat{g}_{t}^{\mathrm{mix}}$ are computed from the same Monte Carlo pool and often partially cancel, improving concentration of $\widehat{r}_{\mathrm{FOC}}$. In implementation, we combine (110) with the inversion stabilizers of Appendix F.5 (e.g. ridge regularization and diagnostic-based fallback to $\pi_{\varphi^{\mathrm{warm}}}$ on unreliable states).
G.2 Interactive distillation: projection-guided training and amortized deployment
Let $\pi_{\varphi}$ be the trainable stage 1 policy network. At intermittent refresh times, freeze a lagged copy $\pi_{\varphi^{-}}$ and run stage 2 under $\pi_{\varphi^{-}}$ to construct a $q$-aggregated projected teacher. This coupling serves two purposes:
  • Training aid. Projection-guided targets can stabilize and accelerate stage 1 optimization once the warm-up policy enters a locally stable regime.
  • Amortization. Stage 2 projection is accurate but Monte-Carlo intensive; distillation compresses it into a single forward pass for deployment.
Teacher definition (computed by stage 2; treated as fixed).
In residual form (110), define the teacher as the $\theta$-blind map
@@ G.2 Interactive distillation: projection-guided training and amortized deployment
where $\widehat{A}_{t}^{\varphi^{-}}$ and $\widehat{g}_{t}^{\mathrm{mix},\varphi^{-}}$ denote the stage 2 estimators computed under the lagged policy $\pi_{\varphi^{-}}$, and $\widehat{r}_{\mathrm{FOC}}^{\varphi^{-}}:=\widehat{A}_{t}^{\varphi^{-}}\pi_{\varphi^{-}}+\widehat{g}_{t}^{\mathrm{mix},\varphi^{-}}$. (We omit stabilizer notation; in code, $(\widehat{A}_{t}^{\varphi^{-}})^{-1}$ is implemented with ridge/diagnostics/skip as needed.)
Student objective (hybrid DPO + distillation penalty).
Train $\pi_{\varphi}$ by combining the ex–ante objective with a proximity penalty to the teacher on the working domain:
$\max_{\varphi}\;J(\varphi)\;-\;\lambda\,\mathbb{E}_{z=(t,x,y)\sim\mu}\Big[\big\|\pi_{\varphi}(z)-\mathrm{stopgrad}\big(\widehat{\pi}^{\mathrm{agg,mix}}_{\varphi^{-}}(z)\big)\big\|^{2}\Big],$ (112)
where $\lambda\geq 0$ controls guidance strength and $\mathrm{stopgrad}(\cdot)$ indicates that gradients are not propagated through stage 2.
Refresh schedule, annealing, and diagnostic gating.
In practice, $\varphi^{-}$ and the teacher are refreshed on a slower timescale than stage 1 updates:
  • Two-time-scale refresh. Hold $\varphi^{-}$ fixed for $n_{\mathrm{refresh}}$ stage 1 steps, then set $\varphi^{-}\leftarrow\varphi$ and recompute the teacher.
  • Anneal $\lambda$. Start with $\lambda=0$ (pure DPO) and increase $\lambda$ only after projection diagnostics on the working domain stabilize.
  • Adaptive teacher selection (gating). Apply teacher loss only on states where projection diagnostics certify reliability (stable $\widehat{A}_{t}$ solve and moderate residual), falling back to pure DPO elsewhere.
Deployment. After training, the deployed controller is the student network $\pi_{\varphi}$, which approximates the stage 2 projected rule with a single forward pass.
Appendix H Stage 2 projection diagnostics
We report Stage 2 diagnostic statistics as a visual supplement to Section 4.2. Each figure uses the same tail-median protocol and layout; see captions for definitions.
@@ Appendix H Stage 2 projection diagnostics
Appendix I Supplementary decomposition diagnostics for Section 5
Tables 45 report Stage 2 decomposition diagnostics at $t=0$ (tail medians over the last six evaluation checkpoints). Diagnostics are reported for the projected (Stage 2 / PG–DPO) rules, since Stage 1 (DPO) and PPO do not expose a compatible myopic/hedging split under our protocol.
@@ Appendix I Supplementary decomposition diagnostics for Section 5
$10^{-3}$ PG–DPO (basic) $7.17\text{\times}{10}^{-06}$ $1.51\text{\times}{10}^{-05}$ $1.60\text{\times}{10}^{-05}$ $1.23\text{\times}{10}^{-05}$
$10^{-3}$ PG–DPO (distill.) $5.20\text{\times}{10}^{-06}$ $9.81\text{\times}{10}^{-06}$ $1.64\text{\times}{10}^{-05}$ $1.62\text{\times}{10}^{-05}$
$10^{-2}$ PG–DPO (basic) $6.21\text{\times}{10}^{-06}$ $1.38\text{\times}{10}^{-05}$ $1.63\text{\times}{10}^{-05}$ $1.42\text{\times}{10}^{-05}$
$10^{-2}$ PG–DPO (distill.) $7.13\text{\times}{10}^{-06}$ $7.13\text{\times}{10}^{-06}$ $1.62\text{\times}{10}^{-05}$ $1.76\text{\times}{10}^{-05}$
$10^{-1}$ PG–DPO (basic) $8.18\text{\times}{10}^{-06}$ $1.41\text{\times}{10}^{-05}$ $3.82\text{\times}{10}^{-05}$ $2.42\text{\times}{10}^{-05}$
$10^{-1}$ PG–DPO (distill.) $6.72\text{\times}{10}^{-06}$ $9.21\text{\times}{10}^{-06}$ $3.67\text{\times}{10}^{-05}$ $3.16\text{\times}{10}^{-05}$
@@
$10^{-3}$ PG–DPO (basic) $1.10\text{\times}{10}^{-05}$ $1.41\text{\times}{10}^{-05}$ $70\text{\times}{10}^{-05} 1.1$$18\text{\times}{10}^{-05}$
$10^{-3}$ PG–DPO (distill.) $7.71\text{\times}{10}^{-06}$ $5.94\text{\times}{10}^{-06}$ $1.84\text{\times}{10}^{-05}$ $1.62\text{\times}{10}^{-05}$
$10^{-2}$ PG–DPO (basic) $7.90\text{\times}{10}^{-06}$ $2.02\text{\times}{10}^{-05}$ $1.46\text{\times}{10}^{-05}$ $1.20\text{\times}{10}^{-05}$
$10^{-2}$ PG–DPO (distill.) $6.00\text{\times}{10}^{-06}$ $1.71\text{\times}{10}^{-05}$ $2.21\text{\times}{10}^{-05}$ $1.43\text{\times}{10}^{-05}$
$10^{-1}$ PG–DPO (basic) $1.24\text{\times}{10}^{-05}$ $1.93\text{\times}{10}^{-04}$ $5.77\text{\times}{10}^{-05}$ $3.14\text{\times}{10}^{-05}$
$10^{-1}$ PG–DPO (distill.) $1.20\text{\times}{10}^{-05}$ $1.90\text{\times}{10}^{-04}$ $7.40\text{\times}{10}^{-05}$ $2.47\text{\times}{10}^{-05}$
Table 4: Myopic-component RMSE at $t=0$ for the projected (Stage 2 / PG–DPO) rules (tail medians).
@@
$10^{-3}$ PG–DPO (basic) $0.994$ $0.988$ $0.991$ $0.990$
$10^{-3}$ PG–DPO (distill.) $0.995$ $0.986$ $0.990$ $0.987$
$10^{-2}$ PG–DPO (basic) $0.993$ $0.989$ $0.992$ $0.988$
$10^{-2}$ PG–DPO (distill.) $0.992$ $0.994$ $0.990$ $0.987$
$10^{-1}$ PG–DPO (basic) $0.988$ $0.990$ $0.936$ $0.932$
$10^{-1}$ PG–DPO (distill.) $0.996$ $0.990$ $0.949$ $0.922$
@@
$10^{-3}$ PG–DPO (basic) $0.988$ $0.988$ $0.993$ $0.990$
$10^{-3}$ PG–DPO (distill.) $0.994$ $0.995$ $0.990$ $0.987$
$10^{-2}$ PG–DPO (basic) $0.994$ $0.976$ $0.992$ $0.988$
$10^{-2}$ PG–DPO (distill.) $0.994$ $0.980$ $0.988$ $0.989$
$10^{-1}$ PG–DPO (basic) $0.990$ $0.005$ $0.668$ $0.851$
$10^{-1}$ PG–DPO (distill.) $0.992$ 0.009-0.009 $0.642$ $0.871$
Table 5: Hedging-direction cosine similarity at $t=0$ for the projected (Stage 2 / PG–DPO) rules (tail medians). Higher is better; negative indicates direction reversal.